# SemanticGen: Video Generation in Semantic Space

### v2

@@ 3 Method
In this section, we present the design of SemanticGen. To generate the video clip $V$ from the text prompt, we leverage pre-trained video diffusion models (Sec. 3.1) and semantic encoders. We start by training a video diffusion model to generate video VAE latents conditioned on their compressed semantic representations (Sec. 3.2). Next, we learn the distribution of the compressed semantic representation from the text input (Sec. 3.3). During inference, we first generate the semantic representation, then map it to the VAE latent space. We also demonstrate that our model can effectively generalize to long video generation (Sec. 3.4). The overview of the model is depicted in Fig. 3.
Figure 3: Overview of SemanticGen. (a) We train a semantic generator to fit the compressed semantic representation distribution of off-the-shelf semantic encoders. (b) We optimized a latent diffusion model for denoising video VAE latents conditioned on their semantic representations. (c) During inference, we integrate the semantic generator and VAE latent generator to achieve high-quality T2V generation.
@@ 3.2 Video Generation with Semantic Embeddings
As a result, we utilize the vision tower of Qwen-2.5-VL [qwen25vl] as our semantic encoder, which is trained with a vision-language alignment objective [clip] on image and video datasets. For video input, it first samples video frames at a lower fps (default 2.0), compresses 14x14 image patches into a single token, and then further compresses along each dimension by a factor of 2. This process ultimately transforms a video $V\in\mathbb{R}^{3\times F\times H\times W}$, into a semantic representation $z_{sem}^{{}^{\prime}}\in\mathbb{R}^{d\times F_{s}/2\times H/28\times W/28}$, where $d$ is the embedding dimension, and $F_{s}$ is the number of video frames sampled as input to the Qwen2.5-VL vision tower. Note that the framework proposed in this paper does not rely on using a specific semantic tokenizer. Other video semantic tokenizers, such as V-JEPA 2 [v-jepa2], VideoMAE 2 [videomae2], and 4DS [4ds], are also compatible with SemanticGen. In this paper, we use Qwen-2.5-VL to validate the effectiveness of the proposed method, and we leave the systematic exploration of other semantic encoders as future work.
@@
The pipeline of the VAE latent generation stage is illustrated in Fig. 3(b). During training, we first feed the input video to the semantic encoder and the learnable MLP to get its compact semantic representation $z_{sem}$, which is then injected into the diffusion model via in-context conditioning [recammaster]. Specifically, we concatenated noised VAE latents $z_{t}$ and compressed semantic representations $z_{sem}$ as the model’s input, i.e., $z_{input}:=[z_{t},z_{sem}]$. To verify that the compressed semantic representation captures the video’s high-level semantics and effectively guides generation, we extract semantic features from a reference video and inject them into the VAE latent generator. The generated video, shown in Fig. 4, preserves the spatial layout and motion patterns of the reference video while differing in fine details. This demonstrates that the compressed semantic representations encode high-level information—such as structure and dynamics, while discarding low-level attributes like texture and color. At the inference stage, $z_{sem}$ is generated by the semantic representation generation model in Sec. 3.3. Similar to RAE [rae], we add noise to $z_{sem}$ to reduce the training-inference gap.
@@ 3.3 Semantic Representation Generation
After training a VAE latent generator to synthesize VAE latents with the compressed semantic representations in Sec. 3.2, we further learn the semantic representation distribution by another video diffusion model. In this stage, we freeze the visual encoder and the MLP, and fine-tune only the latent diffusion model. We observe a significant improvement in convergence speed after regularizing the semantic space with the learnable MLP. The results are summarized in Fig. 8 and Tab. 3. Additionally, we ablate the design of using semantic encoders rather than VAE encoders in Fig. 9, and we observe a significantly faster convergence speed compared to modeling the compressed VAE latents.

### v3

@@ 3 Method
In this section, we present the design of SemanticGen. To generate the video clip $V$ from the text prompt, we leverage pre-trained video diffusion models (Sec. 3.1) and semantic encoders. We start by training a video diffusion model to generate video VAE latents conditioned on their compressed semantic representations (Fig. 3a, Sec. 3.2). Next, we learn the distribution of the compressed semantic representation from the text input (Fig. 3b, Sec. 3.3). During inference, we first generate the semantic representation, then map it to the VAE latent space (Fig. 3c). We also demonstrate that our model can effectively generalize to long video generation (Sec. 3.4).
Figure 3: Overview of SemanticGen. (a) We optimize a latent diffusion model for denoising video VAE latents conditioned on their compressed semantic representations. (b) We train a semantic generator to fit the compressed semantic representation distribution of off-the-shelf semantic encoders. (c) During inference, we integrate the semantic generator and VAE latent generator to achieve high-quality T2V generation. Green: Input; Yellow: Output; Blue: Trainable models; Gray: Frozen models.
@@ 3.2 Video Generation with Semantic Embeddings
As a result, we utilize the vision tower of Qwen-2.5-VL [qwen25vl] as our semantic encoder, which is trained with a vision-language alignment objective [clip] on image and video datasets. For video input, it first samples video frames at a lower fps (default 2.0), compresses 14x14 image patches into a single token, and then further compresses along each dimension by a factor of 2. This process ultimately transforms a video $V\in\mathbb{R}^{3\times F\times H\times W}$, into a semantic representation $z_{sem}^{{}^{\prime}}\in\mathbb{R}^{d\times F_{s}/2\times H/28\times W/28}$, where $d$ is the embedding dimension, and $F_{s}$ is the number of video frames sampled as input to the Qwen2.5-VL vision tower. Note that the framework proposed in this paper does not rely on using a specific semantic tokenizer. Other video semantic tokenizers, such as V-JEPA 2 [v-jepa2], VideoMAE 2 [videomae2], and 4DS [4ds], are also compatible with SemanticGen. In this paper, we use Qwen-2.5-VL to validate the effectiveness of SemanticGen, and we leave the systematic exploration of other semantic encoders as future work.
@@
The pipeline of the VAE latent generation stage is illustrated in Fig. 3a. During training, we first feed the input video to the semantic encoder and the learnable MLP to get its compact semantic representation $z_{sem}$, which is then injected into the diffusion model via in-context conditioning [recammaster]. Specifically, we concatenated noised VAE latents $z_{t}$ and compressed semantic representations $z_{sem}$ as the model’s input, i.e., $z_{input}:=[z_{t},z_{sem}]$. To verify that the compressed semantic representation captures the video’s high-level semantics and effectively guides generation, we extract semantic features from a reference video and inject them into the VAE latent generator. The generated video, shown in Fig. 4, preserves the spatial layout and motion patterns of the reference video while differing in fine details. This demonstrates that the compressed semantic representations encode high-level information—such as structure and dynamics, while discarding low-level attributes like texture and color. At the inference stage, $z_{sem}$ is generated by the semantic representation generation model in Sec. 3.3. Similar to RAE [rae], we add noise to $z_{sem}$ to reduce the training-inference gap.
@@ 3.3 Semantic Representation Generation
After training a VAE latent generator to synthesize VAE latents with the compressed semantic representations in Sec. 3.2 (illustrated in Fig. 3a), we further learn the semantic representation distribution by another video diffusion model (illustrated in Fig. 3b). In this stage, we freeze the visual encoder and the MLP, and fine-tune only the latent diffusion model. We observe a significant improvement in convergence speed after regularizing the semantic space with the learnable MLP. The results are summarized in Fig. 8 and Tab. 3. Additionally, we ablate the design of using semantic encoders rather than VAE encoders in Fig. 9, and we observe a significantly faster convergence speed compared to modeling the compressed VAE latents.